OpenStudy (aonz):

Medal and fan for best answer Halp me please if log A = bt + log P, express A in terms of the other symbols

OpenStudy (anonymous):

Do you mean eliminate the logs from both sides?

OpenStudy (aonz):

no write in terms of A

OpenStudy (aonz):

or make A the subject

OpenStudy (anonymous):

The write it terms of A, just exponentiate each side. Which means, raise each side as a power of e. so \[e^{lnA}=e^{bt+lnP} \] which leaves you with \[A=e^{bt}+P\] If that looks like what you've been doing in class?

OpenStudy (anonymous):

If thats not what you were looking for I apologize and hopefully someone else can hlp

OpenStudy (aonz):

answer at the back of my book sayd A=P10^bt

OpenStudy (aonz):

can try get that answer lol?

OpenStudy (anonymous):

Whoops, thats because I used the natural log. I when its jut log, its base is assumed to be 10. So you do what I did, but with 10 instead of e.

OpenStudy (anonymous):

If you have say, 10^logA , this simplifies to just A. The easiest way to remember logs is "the power to which (the base) must be raised to get whats in the log.

OpenStudy (anonymous):

so, you take 10 raised to the power of each side of your equation. \[10^{logA} = 10^{bt}+10^{logP}\] Which simplified, gives A=10^BT+P

OpenStudy (aonz):

still answer says its times p not +P

OpenStudy (anonymous):

hmm, I guess I fail at logs. http://dl.uncw.edu/digilib/mathematics/algebra/mat111hb/eandl/logprop/logprop.html#sec2 here is what I am referencing. These properties will hopefully help you with whatever Im missing.

OpenStudy (kainui):

If you subtract log P from both sides you get logA-logP which simplifies to log (A/P) and then when you take 10 to both sides you'll get A/P=10^bt and then you can multiply P by both sides to get yours.

OpenStudy (anonymous):

 Now it gives, A= P (10)^bt